the probability of winning is 1/3. Mona plays 6 games. Find probability of winning exactly 5 games.


1
Expert's answer
2022-05-16T23:56:55-0400

We have a Bernoulli trial - exactly two possible outcomes, "success" (Mona wins the game) and "failure" (Mona doesn't win the game) and the probability of success is the same every time the experiment is conducted (Mona plays a game), "p=\cfrac{1}{3}, q=1-\cfrac{1}{3}=\cfrac{2}{3}, n=6."


The probability that Mona wins k games

"P(X=k)=\begin{pmatrix}n\\k\end{pmatrix}\cdot p^k\cdot q^{n-k}=\\
=\begin{pmatrix}6\\k\end{pmatrix}\cdot \begin{pmatrix}\cfrac{1}{3}\end{pmatrix}
^k\cdot \begin{pmatrix}\cfrac{2}{3}\end{pmatrix}^{6-k}=\\
=\cfrac{6!}{k!\cdot(6-k)!}\cdot \begin{pmatrix}\cfrac{1}{3}\end{pmatrix}^k\cdot \begin{pmatrix}\cfrac{2}{3}\end{pmatrix}^{6-k}."

The probability that Mona wins 5 games

"P(X=5)=\cfrac{6!}{5!\cdot1!}\cdot \begin{pmatrix}\cfrac{1}{3}\end{pmatrix}^{5}\cdot \begin{pmatrix}\cfrac{2}{3}\end{pmatrix}^{1}=0.0165."